OpenStudy (anonymous):

air temperature drops approximately 5.5f for each 1,000-foot rise in altitude above earth's surface (up to 30,000ft) a)write a formula that relates temperature t in degrees fahrenheit at altitude h (in thousands of feet) and a ground temperature of 65f. state any restrictions on h. b) find the temperature at 11,000 ft above earth's surface

OpenStudy (anonymous):

air temperature drops approximately 5.5f for each 1,000-foot rise in altitude above earth's surface (up to 30,000ft) a)write a formula that relates temperature t in degrees fahrenheit at altitude h (in thousands of feet) and a ground temperature of 65f. state any restrictions on h. b) find the temperature at 11,000 ft above earth's surface

OpenStudy (cwrw238):

t = 65 - 5.5h h must be <= 30 at 11,000 ft t = 65 - 11*5.5 = 65 - 60.5 = 4.5 degrees fahrenheit